# Large Language Models for Cyber Security: A Systematic Literature Review

### v3

@@ 8.1. Challenges
Model size and deployment. The size of LLMs have seen significant growth over time, escalating from 117M parameters for GPT-1 to 1.5B parameters for GPT-2, and further to 175B parameters for GPT-3 (Yang et al., 2023b). Models with billions or even trillions of parameters present substantial challenges in terms of storage, memory, and computational demands (Fedus et al., 2022). This can potentially impede the deployment of LLMs, particularly in scenarios where developers lack access to potent GPUs or TPUs, especially in resource-constrained environments necessitating real-time deployment. CodeBERT (Feng et al., 2020) emerged in 2019 as a pre-trained model featuring 125M parameters and a model size of 476MB. Recent models like Codex (Chen et al., 2021) and CodeGen (Nijkamp et al., 2023) have surpassed 100 billion parameters, with model sizes exceeding 100GB. Larger sizes entail more computational resources and higher time costs. For instance, training the GPT-Neox-20B model (Black et al., 2021) mandates 825GB of raw text data and deployment on 8 NVIDIA A100-SXM4-40GB graphics processing units (GPUs). Each GPU comes with a price tag of over \$6,000, and the training duration spans 1,830 hours or roughly 76 days. These instances underscore the substantial computational costs linked with training LLMs. Additionally, these platforms entail notable energy expenses, with LLM-based platforms projected to markedly amplify energy consumption (Rillig et al., 2023). Some vendors like OpenAI and Google provide online APIs for LLMs to alleviate user usage costs, while researchers explore methods to curtail LLM scale. Hsieh et al. (Hsieh et al., 2023) proposed step-by-step distillation to diminish the data and model size necessary for LLM training, with their findings showcasing that a T5 model with only 770MB surpassed a 540B PaLM.
@@
Data scarcity. In Section 6, we conducted an extensive examination of the datasets and data preprocessing procedures employed in the 118 studies. Our analysis unveiled the heavy reliance of LLMs on a diverse array of datasets for training and fine-tuning. The findings underscore the challenge of data scarcity encountered by LLMs when tackling security tasks. The quality, diversity, and volume of data directly influence the performance and generalization capabilities of these models. Given their scale, LLMs typically necessitate substantial data volumes to capture nuanced distinctions, yet acquiring such data poses significant challenges. Many specific security tasks suffer from a dearth of high-quality and robust publicly available datasets. Relying on limited or biased datasets may result in models inheriting these biases, leading to skewed or inaccurate predictions. Furthermore, there is a concern regarding the risk of benchmark data contamination, where existing research may involve redundant filtering of native data, potentially resulting in overlap between training and testing datasets, thus inflating performance metrics (Lee et al., 2022). Additionally, we raise serious apprehensions regarding the inclusion of personally private information, such as phone numbers and email addresses, in training corpora when LLMs are employed for information and content security tasks, which precipitate privacy breaches during the prompting process (El-Mhamdi et al., 2023).
@@
The generalization capability of LLMs pertains to their ability to consistently and accurately execute tasks across diverse tasks, datasets, or domains beyond their training environment. Despite undergoing extensive pre-training on large datasets to acquire broad knowledge, the absence of specialized expertise can present challenges when LLMs encounter tasks beyond their pre-training scope, especially in the cybersecurity domain. As discussed in Section  3, we explored the utilization of LLMs in 21 security tasks spanning five security domains. We observed substantial variations in the context and semantics of code or documents across different domains and task specifications. To ensure LLMs demonstrate robust generalization, meticulous fine-tuning, validation, and continuous feedback loops on datasets from various security tasks are imperative. Without these measures, there’s a risk of models overfitting to their training data, thus limiting their efficacy in diverse real-world scenarios.
@@
Ensuring interpretability and trustworthiness is paramount when integrating LLMs into security tasks, particularly given the sensitive nature of security requirements and the need for rigorous scrutiny of model outputs. The challenge lies in comprehending how these models make decisions, as the black-box nature of LLMs often impedes explanations for why or how specific outputs or recommendations are generated for security needs. Recent research (Perry et al., 2023; Wang et al., 2023d) has underscored that artificial intelligence-generated content (AIGC) introduces additional security risks, including privacy breaches, dissemination of forged information, and the generation of vulnerable code. The absence of interpretability and trustworthiness can breed user uncertainty and reluctance, as stakeholders may hesitate to rely on LLMs for security tasks without a clear understanding of their decision-making process or adherence to security requirements. Establishing trust in LLMs necessitates the development of technologies and tools that offer deeper insights into model internals, empowering developers to comprehend the rationale behind generated outputs. Improving interpretability and trustworthiness can ultimately foster the widespread adoption of cost-effective automation in the cybersecurity domain, fostering more efficient and effective security practices. Many LLMs lack open-source availability, and questions persist regarding the data on which they were trained, as well as the quality, sources, and ownership of the training data, raising concerns about ownership regarding LLM-generated tasks. Moreover, there is the looming threat of various adversarial attacks, including tactics to guide LLMs to circumvent security measures and expose their original training data (Deng et al., 2024).
@@ 8.2. Opportunities
Inter-model interaction of LLMs. Our examination indicates that LLMs have progressed significantly in tackling various security challenges. However, as security tasks become more complex, there’s a need for more sophisticated and tailored solutions. As outlined in Section 5, one promising avenue is collaborative model interaction through external augmentation methods. This approach involves integrating multiple LLMs (Yang et al., 2023a) or combining LLMs with specialized machine learning models (Tang et al., 2023; Cai et al., 2024) to improve task efficiency while simplifying complex steps. By harnessing the strengths of different models collectively, we anticipate that LLMs can deliver more precise and higher-quality outcomes for intricate security tasks.
Impact and applications of ChatGPT. In recent academic research, ChatGPT has garnered considerable attention, appearing in over half of the 127 papers we analyzed. It has been utilized to tackle specific security tasks, highlighting its growing influence and acceptance in academia. Researchers have favored ChatGPT due to its computational efficiency, versatility across tasks, and potential cost-effectiveness compared to other LLMs and LLM-based applications (Laskar et al., 2023b). Beyond generating task solutions, ChatGPT promotes collaboration, signaling a broader effort to integrate advanced natural language understanding into traditional cybersecurity practices (Deng et al., 2023a; Qammar et al., 2023). By closely examining these trends, we can anticipate pathways for LLMs and applications like ChatGPT to contribute to more robust, efficient, and collaborative cybersecurity solutions. These insights highlight the transformative potential of LLMs in shaping the future cybersecurity landscape.
@@
External retrieval and tools for LLM. LLMs have demonstrated impressive performance across diverse security tasks, but they are not immune to inherent limitations, including a lack of domain expertise (Kandpal et al., 2023), a tendency to generate hallucinations (Zhang et al., 2023e), weak mathematical capabilities, and a lack of interpretability. Therefore, a feasible approach to enhancing their capabilities is to enable them to interact with the external world, acquiring knowledge in various forms and manners to improve the factualness and rationality of generated security task solutions. One viable solution is to provide external knowledge bases for LLMs, augmenting content generation with retrieval-based methods to retrieve task-relevant data for LLM outputs (Gao et al., 2024; Du and Yu, 2023). Another approach is to incorporate external specialized tools to provide real-time interactive feedback to guide LLMs (Alrashedy and Aljasser, 2024; Armengol-Estapé et al., 2024), combining the results of specialized analytical tools to steer LLMs towards robust and consistent security task solutions. We believe that incorporating external retrieval and tools is a competitive choice for improving the performance of LLM4Security.
Addressing challenges in specific domains. Numerous cybersecurity domains, such as network security and hardware security, encounter a dearth of open-source datasets, impeding the integration of LLMs into these specialized fields (Sun et al., 2023b). Future endeavors may prioritize the development of domain-specific datasets and the refinement of LLMs to address the distinctive challenges and nuances within these domains. Collaborating with domain experts and practitioners is crucial for gathering relevant data, and fine-tuning LLMs with this data can improve their effectiveness and alignment with each domain’s specific requirements. This collaborative approach helps LLMs address real-world challenges across different cybersecurity domains (Bubeck et al., 2023).
@@
Integrating new input formats. In our research, we noticed that LLMs in security tasks typically use input formats from code-based and text-based datasets. The introduction of new input formats based on natural language, like voice and images, as well as multimodal inputs such as video demonstrations, presents an opportunity to enhance LLMs’ ability to understand and process various user needs (Yin et al., 2024). Integrating speech can improve user-model interaction, allowing for more natural and context-rich communication. Images can visually represent security task processes and requirements, providing LLMs with additional perspectives. Moreover, multimodal inputs combining text, audio, and visuals can offer a more comprehensive contextual understanding, leading to more accurate and contextually relevant security solutions.
@@ 8.3. Roadmap
Incorporating security knowledge into LLMs. A key direction for the future is to integrate specialized security task solutions and knowledge from the cybersecurity domain into LLMs to overcome potential hallucinations and errors (Aghaei and Al-Shaer, 2023; Ling et al., 2023). This integration aims to enhance LLMs’ ability to address security tasks, especially those requiring a significant amount of domain expertise, such as penetration testing (Deng et al., 2023a; Temara, 2023), hardware vulnerability detection (Dessouky et al., 2019), log analysis (Liu et al., 2023b; Karlsen et al., 2023), and more. Embedding rules and best practices from specific security domains into these models will better represent task requirements, enabling LLMs to generate robust and consistent security task solutions.
Security agent: integrating external augmentation and LLMs. We have witnessed the unprecedented potential of applying LLMs to solve security tasks, almost overturning traditional security task solutions in LLM4Security (Liang et al., 2024; Chen et al., 2023c; Xia et al., 2024). However, the inherent lack of domain-specific knowledge and hallucinations in LLMs restrict their ability to perceive task requirements or environments with high quality (Zhang et al., 2023e). AI Agents are artificial entities that perceive the environment, make decisions, and take actions. Currently, they are considered the most promising tool for achieving the pursuit of achieving or surpassing human-level intelligence in specific domains (Xi et al., 2023). We summarized the external enhancement techniques introduced in LLM4Security in Section 5, optimizing LLMs’ performance in security tasks across multiple dimensions, including input, model, and output (Jin et al., 2023; Perrina et al., 2023; Chen et al., 2023c). Security operators can specify specific external enhancement strategies for security tasks and integrate them with LLMs to achieve automated security AI agents with continuous interaction within the system.
@@
Security for Large Language Models (Security4LLM). LLMs have gained considerable traction in the security sector, showcasing their potential in security-related endeavors. Nonetheless, delving into the internal security assessment of LLMs remains a pressing area for investigation (Yao et al., 2024). The intricate nature of LLMs renders them vulnerable to attacks, necessitating innovative strategies to fortify the models themselves (Deng et al., 2024; Liu et al., 2023a; Glukhov et al., 2023). Previous studies have identified vulnerabilities in LLMs like jailbreaking and malicious prompt injection, resulting in the exposure of model training data or sensitive user chat records (Liu et al., 2023a; Derner and Batistič, 2023; Greshake et al., 2023). Considering that the inputs for security tasks often involve security-sensitive data (such as system logs and vulnerability code in programs) (Qi et al., 2023a; Pearce et al., 2023), the leakage of such information would pose significant cybersecurity risks. An intriguing avenue for future research is to empower LLMs to autonomously detect and identify their vulnerabilities. Specifically, efforts could focus on enabling LLMs to generate patches for their underlying code, thus bolstering their inherent security, rather than solely implementing program restrictions at the user interaction layer. Given this scenario, future research should adopt a balanced approach, striving to utilize LLMs for automating cost-effective completion of security tasks while simultaneously developing techniques to safeguard the LLMs themselves. This dual focus is pivotal for fully harnessing the potential of LLMs in enhancing cybersecurity and ensuring compliance with cyber systems.

### v4

@@ 8.1. Challenges
Model size and deployment. The size of LLMs have seen significant growth over time, escalating from 117M parameters for GPT-1 to 1.5B parameters for GPT-2, and further to 175B parameters for GPT-3 (Yang et al., 2023a). Models with billions or even trillions of parameters present substantial challenges in terms of storage, memory, and computational demands (Fedus et al., 2022). This can potentially impede the deployment of LLMs, particularly in scenarios where developers lack access to potent GPUs or TPUs, especially in resource-constrained environments necessitating real-time deployment. CodeBERT (Feng et al., 2020) emerged in 2019 as a pre-trained model featuring 125M parameters and a model size of 476MB. Recent models like Codex (Chen et al., 2021) and CodeGen (Nijkamp et al., 2023) have surpassed 100 billion parameters, with model sizes exceeding 100GB. Larger sizes entail more computational resources and higher time costs. For instance, training the GPT-Neox-20B model (Black et al., 2021) mandates 825GB of raw text data and deployment on 8 NVIDIA A100-SXM4-40GB graphics processing units (GPUs). Each GPU comes with a price tag of over \$6,000, and the training duration spans 1,830 hours or roughly 76 days. These instances underscore the substantial computational costs linked with training LLMs. Additionally, these platforms entail notable energy expenses, with LLM-based platforms projected to markedly amplify energy consumption (Rillig et al., 2023). Some vendors like OpenAI and Google provide online APIs for LLMs to alleviate user usage costs, while researchers explore methods to curtail LLM scale. Hsieh et al. (Hsieh et al., 2023) proposed step-by-step distillation to diminish the data and model size necessary for LLM training, with their findings showcasing that a T5 model with only 770MB surpassed a 540B PaLM.
@@
Data Quality, Availability, and Security Challenges. In Section 6, we conducted an extensive examination of the datasets and data preprocessing procedures employed in the 118 studies. Our analysis unveiled the heavy reliance of LLMs on a diverse array of datasets for training and fine-tuning. The findings underscore the challenge of data scarcity encountered by LLMs when tackling security tasks. The quality, diversity, and volume of data directly influence the performance and generalization capabilities of these models. Given their scale, LLMs typically necessitate substantial data volumes to capture nuanced distinctions, yet acquiring such data poses significant challenges. Many specific security tasks suffer from a dearth of high-quality and robust publicly available datasets. Relying on limited or biased datasets may result in models inheriting these biases, leading to skewed or inaccurate predictions. Furthermore, there is a concern regarding the risk of benchmark data contamination, where existing research may involve redundant filtering of native data, potentially resulting in overlap between training and testing datasets, thus inflating performance metrics (Lee et al., 2022). Additionally, we raise serious apprehensions regarding the inclusion of personally private information, such as phone numbers and email addresses, in training corpora when LLMs are employed for information and content security tasks, which precipitate privacy breaches during the prompting process (El-Mhamdi et al., 2023).
@@
The generalization capability of LLMs pertains to their ability to consistently and accurately execute tasks across diverse tasks, datasets, or domains beyond their training environment. Despite undergoing extensive pre-training on large datasets to acquire broad knowledge, the absence of specialized expertise can present challenges when LLMs encounter tasks beyond their pre-training scope, especially in the cybersecurity domain. As discussed in Section  4, we explored the utilization of LLMs in 21 security tasks spanning five security domains. We observed substantial variations in the context and semantics of code or documents across different domains and task specifications. To ensure LLMs demonstrate robust generalization, meticulous fine-tuning, validation, and continuous feedback loops on datasets from various security tasks are imperative. Without these measures, there’s a risk of models overfitting to their training data, thus limiting their efficacy in diverse real-world scenarios.
@@
Ensuring interpretability and trustworthiness is paramount when integrating LLMs into security tasks, particularly given the sensitive nature of security requirements and the need for rigorous scrutiny of model outputs. The challenge lies in comprehending how these models make decisions, as the black-box nature of LLMs often impedes explanations for why or how specific outputs or recommendations are generated for security needs. While LLM self-explanations are often post-hoc rationalizations rather than reflections of internal workings (Freiberger et al., 2025; Sarkar, 2024), research explores methods to enhance their transparency and utility in security contexts. Techniques include promoting step-by-step reasoning during generation (e.g., Chain-of-Thought) which can improve output quality (Kojima et al., 2023; Freiberger et al., 2025), grounding explanations in source data using RAG or citation mechanisms, visualizing attention weights as proxies (El Houda Dehimi and Tolba, 2024), defining structured evaluation criteria to constrain the LLM’s reasoning process, and applying established XAI methods like LIME and SHAP to explain the behavior of models used within security workflows (Lim et al., 2025). Furthermore, assessing the reliability of LLM judgments can be aided by estimating output uncertainty or checking for consistency across multiple generated responses (Varshney et al., 2023). A promising approach involves using LLMs specifically as ’explanation translators,’ taking technical outputs from XAI tools (like feature importance scores from LIME/SHAP) and converting them into accessible, natural language descriptions tailored for end-users, as demonstrated in phishing detection (Lim et al., 2025).
@@
Recent research (Perry et al., 2023; Wang et al., 2023e) has also underscored that artificial intelligence-generated content (AIGC) introduces additional security risks, including privacy breaches, dissemination of forged information, and the generation of vulnerable code. The absence of interpretability and trustworthiness can breed user uncertainty and reluctance, as stakeholders may hesitate to rely on LLMs for security tasks without a clear understanding of their decision-making process or adherence to security requirements. Establishing trust in LLMs necessitates further development and application of technologies and tools that offer deeper insights into model internals and output rationale, empowering developers and users to comprehend and verify the reasoning (Freiberger et al., 2025). Improving interpretability and trustworthiness through such methods can ultimately foster the widespread adoption of cost-effective automation in the cybersecurity domain, fostering more efficient and effective security practices. However, many LLMs lack open-source availability, and questions persist regarding the data on which they were trained, as well as the quality, sources, and ownership of the training data, raising concerns about ownership regarding LLM-generated tasks. Moreover, there is the looming threat of various adversarial attacks, including tactics to guide LLMs to circumvent security measures and expose their original training data (Deng et al., 2024a).
8.1.4. Limitations in Modern Cryptanalysis
A significant challenge to the broad applicability of LLMs is evident in the field of modern cryptanalysis. Current LLMs operate as probabilistic models, proficient at identifying statistical patterns within extensive textual datasets (Gu et al., 2024). This operational paradigm inherently conflicts with the deterministic, precise, and mathematically rigorous demands of cryptanalysis, which typically necessitates profound insights into areas like number theory and algebra to devise novel algorithms for computationally hard problems (Swenson, 2008; Bogdanov et al., 2016). LLMs demonstrate notable limitations in executing complex multi-step mathematical computations, engaging in formal logical deduction, and managing abstract conceptualizations, often leading to errors or diminished performance when faced with irrelevant contextual information (Ahn et al., 2024; Havrilla and Iyer, 2024). Moreover, the data-centric learning approach of LLMs is fundamentally misaligned with the objective of discovering new attack methodologies for contemporary ciphers, primarily because datasets illustrating successful exploits against such robust systems are, by design, unavailable for training. The tendency of LLMs to generate specious yet plausible information (i.e., hallucinations) further compromises their utility in cryptanalytic applications where exactitude is non-negotiable (Zhang et al., 2023d). Thus, the core architecture of LLMs as sophisticated language processors, rather than as deterministic logical inference engines, establishes a fundamental incompatibility with the stringent requirements for breaking modern cryptographic systems.
@@ 8.2. Opportunities
Inter-model interaction of LLMs. Our examination indicates that LLMs have progressed significantly in tackling various security challenges. However, as security tasks become more complex, there’s a need for more sophisticated and tailored solutions. As outlined in Section 5, one promising avenue is collaborative model interaction through external augmentation methods. This approach involves integrating multiple LLMs (Yang et al., 2024a) or combining LLMs with specialized machine learning models (Tang et al., 2023; Cai et al., 2024a) to improve task efficiency while simplifying complex steps. By harnessing the strengths of different models collectively, we anticipate that LLMs can deliver more precise and higher-quality outcomes for intricate security tasks.
Impact and applications of ChatGPT. In recent academic research, ChatGPT has garnered considerable attention, appearing in over half of the 185 papers we analyzed. It has been utilized to tackle specific security tasks, highlighting its growing influence and acceptance in academia. Researchers have favored ChatGPT due to its computational efficiency, versatility across tasks, and potential cost-effectiveness compared to other LLMs and LLM-based applications (Laskar et al., 2023b). Beyond generating task solutions, ChatGPT promotes collaboration, signaling a broader effort to integrate advanced natural language understanding into traditional cybersecurity practices (Deng et al., 2024b; Qammar et al., 2023). By closely examining these trends, we can anticipate pathways for LLMs and applications like ChatGPT to contribute to more robust, efficient, and collaborative cybersecurity solutions. These insights highlight the transformative potential of LLMs in shaping the future cybersecurity landscape.
@@
External retrieval and tools for LLM. LLMs have demonstrated impressive performance across diverse security tasks, but they are not immune to inherent limitations, including a lack of domain expertise (Kandpal et al., 2023), a tendency to generate hallucinations (Zhang et al., 2023d), weak mathematical capabilities, and a lack of interpretability. Therefore, a feasible approach to enhancing their capabilities is to enable them to interact with the external world, acquiring knowledge in various forms and manners to improve the factualness and rationality of generated security task solutions. One viable solution is to provide external knowledge bases for LLMs, augmenting content generation with retrieval-based methods to retrieve task-relevant data for LLM outputs (Gao et al., 2024; Du and Yu, 2023). Another approach is to incorporate external specialized tools to provide real-time interactive feedback to guide LLMs (Alrashedy and Aljasser, 2024; Armengol-Estapé et al., 2024), combining the results of specialized analytical tools to steer LLMs towards robust and consistent security task solutions. We believe that incorporating external retrieval and tools is a competitive choice for improving the performance of LLM4Security.
Addressing challenges in specific domains. Numerous cybersecurity domains, such as network security and hardware security, encounter a dearth of open-source datasets, impeding the integration of LLMs into these specialized fields (Sun et al., 2024). Future endeavors may prioritize the development of domain-specific datasets and the refinement of LLMs to address the distinctive challenges and nuances within these domains. Collaborating with domain experts and practitioners is crucial for gathering relevant data, and fine-tuning LLMs with this data can improve their effectiveness and alignment with each domain’s specific requirements. This collaborative approach helps LLMs address real-world challenges across different cybersecurity domains (Bubeck et al., 2023).
@@
Integrating new input formats. In our research, we noticed that LLMs in security tasks typically use input formats from code-based and text-based datasets. The introduction of new input formats based on natural language, like voice and images, as well as multimodal inputs such as video demonstrations, presents an opportunity to enhance LLMs’ ability to understand and process various user needs (Yin et al., 2024a). Integrating speech can improve user-model interaction, allowing for more natural and context-rich communication. Images can visually represent security task processes and requirements, providing LLMs with additional perspectives. Moreover, multimodal inputs combining text, audio, and visuals can offer a more comprehensive contextual understanding, leading to more accurate and contextually relevant security solutions.
@@ 8.3. Roadmap
Incorporating security knowledge into LLMs. A key direction for the future is to integrate specialized security task solutions and knowledge from the cybersecurity domain into LLMs to overcome potential hallucinations and errors (Aghaei and Al-Shaer, 2023; Ling et al., 2023). This integration aims to enhance LLMs’ ability to address security tasks, especially those requiring a significant amount of domain expertise, such as penetration testing (Deng et al., 2024b; Temara, 2023), hardware vulnerability detection (Dessouky et al., 2019), log analysis (Liu et al., 2023b; Karlsen et al., 2023), and more. Embedding rules and best practices from specific security domains into these models will better represent task requirements, enabling LLMs to generate robust and consistent security task solutions.
Security agent: integrating external augmentation and LLMs. We have witnessed the unprecedented potential of applying LLMs to solve security tasks, almost overturning traditional security task solutions in LLM4Security (Liang et al., 2024; Chen et al., 2023c; Xia et al., 2024). However, the inherent lack of domain-specific knowledge and hallucinations in LLMs restrict their ability to perceive task requirements or environments with high quality (Zhang et al., 2023d). AI Agents are artificial entities that perceive the environment, make decisions, and take actions. Currently, they are considered the most promising tool for achieving the pursuit of achieving or surpassing human-level intelligence in specific domains (Xi et al., 2023). We summarized the external enhancement techniques introduced in LLM4Security in Section 5, optimizing LLMs’ performance in security tasks across multiple dimensions, including input, model, and output (Jin et al., 2023; Perrina et al., 2023; Chen et al., 2023c). Security operators can specify specific external enhancement strategies for security tasks and integrate them with LLMs to achieve automated security AI agents with continuous interaction within the system.
@@
Security for Large Language Models (Security4LLM). LLMs have gained considerable traction in the security sector, showcasing their potential in security-related endeavors. Nonetheless, delving into the internal security assessment of LLMs themselves remains a pressing area for investigation (Yao et al., 2024). The intricate nature of LLMs renders them vulnerable to a wide range of attacks, necessitating innovative strategies to fortify the models (Deng et al., 2024a; Liu et al., 2023a; Glukhov et al., 2023; Yao et al., 2024). These vulnerabilities can be broadly categorized (Yao et al., 2024). AI-inherent vulnerabilities stem from the machine learning aspects and include adversarial attacks like data poisoning (manipulating training data to compromise the model (Wan et al., 2023)) and backdoor attacks (embedding hidden triggers), as well as privacy-compromising inference attacks (attribute inference (Staab et al., 2024), membership inference (Mireshghallah et al., 2022)) and extraction attacks (recovering sensitive training data (Carlini et al., 2021)). Instruction tuning attacks, such as jailbreaking (bypassing safety restrictions (Deng et al., 2024a)) and prompt injection (manipulating prompts to elicit unintended behavior (Greshake et al., 2023)), exploit the way LLMs process instructions. Non-AI inherent vulnerabilities may arise from the surrounding infrastructure, such as remote code execution (RCE) in integrated applications (Liu et al., 2024b) or supply chain vulnerabilities involving third-party plugins (Iqbal et al., 2025).
Furthermore, LLMs can themselves be misused as tools for offensive purposes (Yao et al., 2024)). Their advanced generation capabilities can be exploited to create sophisticated phishing emails (Heiding et al., 2024b), generate malware (Botacin, 2023b), spread misinformation (Chen and Shu, 2024), or facilitate social engineering attacks (Falade, 2023). User-level attacks are particularly prevalent due to the human-like reasoning abilities of LLMs (Yao et al., 2024). Considering that the inputs and outputs for security tasks often involve security-sensitive data (such as system logs or vulnerability code in programs) (Qi et al., 2023a; Pearce et al., 2023), the potential for misuse or data leakage poses significant cybersecurity risks.
Mitigating these threats requires a multi-faceted approach (Yao et al., 2024). Defense strategies can be implemented at different stages: improving model architecture (e.g., considering capacity or sparsity (Li et al., 2022)), during training (e.g., corpora cleaning (Subramani et al., 2023), adversarial training (Wang et al., 2019), safe instruction tuning (Bianchi et al., 2024), applying differential privacy (Li et al., 2022)), and during inference (e.g., instruction pre-processing (Jain et al., 2023), in-process malicious use detection (Wang et al., 2023a), and output post-processing like self-critique (Kadavath et al., 2022)). An intriguing avenue for future research is to empower LLMs to autonomously detect and identify their own vulnerabilities. Specifically, efforts could focus on enabling LLMs to generate patches for their underlying code, thus bolstering their inherent security, rather than solely implementing program restrictions at the user interaction layer. Given this scenario, future research should adopt a balanced approach, striving to utilize LLMs for automating cost-effective completion of security tasks while simultaneously developing and evaluating robust techniques to safeguard the LLMs themselves. This dual focus is pivotal for fully harnessing the potential of LLMs in enhancing cybersecurity and ensuring compliance with cyber systems.